Two dice are rolled. What are the number of ways an 8 can appear from the throw of the dice?
Found 2 solutions by Nate, funmath:
Answer by Nate(3500)

There is a total of five ways.
6 : 2
2 : 6
5 : 3
3 : 5
4 : 4
This tells you the number of possibilities:
f(x) = -|x - 7| + 6
f(8) = -|8 - 7| + 6 = -|1| + 6 = -1 + 6 = 5

Answer by funmath(2933)

Two dice are rolled. What are the number of ways an 8 can appear from the throw of the dice?
A die has 1-6 dots possible.
Below you see the ways you can roll 8 with 2 dice.
(2,6),(3,5),(4,4),(5,3),(6,2)
There are 5 ways to roll an two dice with the sum of 8.